OpenStudy (anonymous):

Point R is at (2, 1.2) and Point T is at (2, 2.5) on a coordinate grid. The distance between the two points is ____. (Input numbers and decimal point only, such as 8.2.)

pooja195 (pooja195):

You need the distance formula \[\huge~\rm~\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

pooja195 (pooja195):

Then you have to do the math :)

pooja195 (pooja195):

x1= 2 x2=2 y1=1.2 y2=2.5 \[\huge~\rm~\sqrt{(2-2)^2+(2.5-1.2)^2}\]

pooja195 (pooja195):

Solve

OpenStudy (anonymous):

i still dont understand

pooja195 (pooja195):

Ok are you fimiliar with the distance formula?

OpenStudy (anonymous):

nope

pooja195 (pooja195):

the distance formula \[\huge~\rm~\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] We are trying to figure out the distance between the 2 point right? This formula will help us find the distance understood?

OpenStudy (anonymous):

ok

pooja195 (pooja195):

So you do know that a coordinate is set up like this right? (x,y) <---

OpenStudy (anonymous):

yea

pooja195 (pooja195):

Ok next we need know that this is how we determine what goes where.... (x1,y1)(x2,y2) understood?

pooja195 (pooja195):

@nomymalik

pooja195 (pooja195):

x1= 2 x2=2 y1=1.2 y2=2.5 \[\huge~\rm~\sqrt{(2-2)^2+(2.5-1.2)^2}\] Here i have just used the numbers given in the directions ad substituted them into the formulathat i showed you earlier (distance formula)

pooja195 (pooja195):

From here you can solve it using a calculator or by hand whatever floats your boat :P let me know what you get :)